My car get 30mpg and my husbands get 25mpg, if we both traveled 200 miles per week and the price of gas is $4.00 a gallon, what is the difference in savings
Answer by flame8855(424)

200 miles / 30 = 20/3 gallons
200/25= 8 gallons
difference = 4(8-20/3)= 16/3 dollars